Suppose in a random sample of 750 people, 135 of them smoke regularly. Find the 94% confidence interval for the population proportion of smokers. Interpret this interval.
1
Expert's answer
2020-03-31T15:28:16-0400

We need to construct the 94% confidence interval for the population proportion. We have been provided with the following information about the number of favorable cases:

Favorable classes "X=135"

Sample size "N=750"

The sample proportion is computed as follows, based on the sample size "N=750" and the number of favorable cases "X=135:"


"\hat{p}={X \over N}={135 \over 750}=0.18"

The critical value for "\alpha=0.06" is "z_c=z_{1-\alpha/2}=1.881." The corresponding confidence interval is computed as shown below:


"CI(Proportion)=\\\big(\hat{p}-z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}}, \hat{p}+z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}}\big)="

"=\big(0.18-1.881\sqrt{{0.18(1-0.18) \over 750}}, 0.18+1.881\sqrt{{0.18(1-0.18) \over 750}}\big)="

"=(0.1536,0.2064)"

Therefore, based on the data provided, the 94% confidence interval for the population proportion is "(0.1536,0.2064)," which indicates that we are 94% confident that the true population proportion "p" lies in the interval "(0.1536,0.2064)."